OpenStudy (anonymous):

What does it mean for a number to be the "solution" to an equation? Give an example of a variable equation and its solution

OpenStudy (anonymous):

Well like x+3=0 and -3 would be the solution

OpenStudy (anonymous):

let the eqn be |x|=2 solution's are -2,2 so it has two solution's the number values of x is known as no of solutions x^2-3x+2=0 is an another eqn here x has two values that is (x-1)(x-2)=0 so x can be x= 1 or x= 2 so we have 2 solution

OpenStudy (anonymous):

To find a "solution" of an equation usually means to find the unknown variable (usually x, a, b, c, etc). You will usually be given with 1 or 2 equations with 1 or 2 unknown variables. So for example, if it says to find a solution for this equation: 2x + 4 = 6 it means to find the x so it would be 2x = 6-4 2x = 2 x = 2 2 would be the solution for x. at times, you'll be given 2 equations with 2 variables and will be asked to solve for the solution for ex: if you were given 2x + 2 = 4 and 2b = x then first, solve for x. 2x = 4-2 2x = 2 x = 1 then solve for b by plugging in x into that other equation. 2b = 1 b = 1/2